OpenStudy (anonymous):

Given that P = (-4, 11) and Q = (-5, 8), find the component form and magnitude of QP @iPwnBunnies

OpenStudy (anonymous):

I know it is <1, 3>

OpenStudy (anonymous):

but would the mag be 10 or sqrt of 10?

OpenStudy (anonymous):

@iPwnBunnies please!!

OpenStudy (ipwnbunnies):

Magnitude would be derived from the Pythagorean Theorem \[\sqrt{1^2 + 3^2} = \sqrt{10}\]

OpenStudy (anonymous):

thanks!